russianmafiya:

Can you factor 3x(x−2)(x+8

umm:

Given: 3x(x−2)(x+8) =(3x(x−2))(x+8) =(3x(x−2))(x)+(3x(x−2))(8) =3x^3−6x^2+24x^2−48x =3x^3+18x^2−48x Therefore your answer is: =3x^3+18x^2−48x